OpenStudy (anonymous):

Find the min of g(t)=abs(t-2) + 10 on the interval [-10,10]

hero (hero):

saifoo got this

OpenStudy (saifoo.khan):

|t - 2| + 10 Vertex = (2, 10)

OpenStudy (saifoo.khan):

Min = 10

OpenStudy (anonymous):

show work?

OpenStudy (saifoo.khan):

How can i show my mental work? D:

hero (hero):

That is his idea of showing work, lol

OpenStudy (anonymous):

lols

OpenStudy (anonymous):

2-2= 0 +10 = 10?

OpenStudy (saifoo.khan):

t - 2 = 0 t = 2

OpenStudy (anonymous):

what about the max of g(t)=l t-6 l + 7 on the interval [-10,10]

OpenStudy (anonymous):

im getting 10 again

hero (hero):

You shouldn't

OpenStudy (saifoo.khan):

Try to learn Mario, t - 6 = 0 t = 6 (6, 7)

hero (hero):

It says max, not min

hero (hero):

I'm sticking with 23 as the max

OpenStudy (anonymous):

its 23. show me how u got that please.

hero (hero):

substitute t = -10

OpenStudy (anonymous):

ok i see